The following data represent the social ambivalence scores for 15 people as measured by a psychological test. (The higher the score, the stronger the ambivalence.) 8 12 11 15 14 10 8 3 8 7 21 12 9 19 11 (a) Guess the value of s using the range approximation. s ≈ (b) Calculate x for the 15 social ambivalence scores. Calculate s for the 15 social ambivalence scores. (c) What fraction of the scores actually lie in the interval x ± 2s? (Round your answer to two decimal places.).
Answer #1

S.NO (X-2)* 10.24 WN = wc O (X-) -3.20 0.80 -0.20 3.80 2.80 -1.20 -3.20 -8.20 -3.20 -4.20 9.80 0.80 -2.20 7.80 -0.20 ONS 8888

a)

from above approximate s =range/4 =(max-min)/4 =(21-3)/4 =4.5

b)

xbar =11.20

s =4.65

d)

range of values between x-/+ 2s =(11.20 -/+ 2*4.65) = (1.90 to 20.50)

therefore t fraction of the scores actually lie in the interval x ± 2s =P(1.90 <X<20.50)=(14/15)=0.93